OpenStudy (anonymous):

Help Please (i^349)^-i^456 Help Please (i^349)^-i^456 @Mathematics

OpenStudy (anonymous):

kind of confused with the equation can u please put brackets to clarify

OpenStudy (anonymous):

I don't know what you mean by putting brackets, there are brackets :) I can write out the problem and take a picture of it

OpenStudy (anonymous):

Well i am getting confused as to the powers

OpenStudy (anonymous):

The last part of the equation is confusing me

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

-1 -1 = -2

OpenStudy (anonymous):

Ishaan how did you get that?

OpenStudy (anonymous):

oh lol sorry I didn't see the complete question sorry \[\Huge{i^{349^{-i^{456}}}}\] It's like this Right?

OpenStudy (anonymous):

yep just like that

OpenStudy (nocipher):

Note that 349 is congruent to 1 modulo 4 and 456 is congruent to 0 modulo 4. Then i^349=i and -i^456=-1 so i^(-1) = -i since 1/i = -i.

OpenStudy (anonymous):

\[\large i^{456} = 1\]

OpenStudy (anonymous):

i^349 = i i^ (-1) yeah what nocipher said

OpenStudy (anonymous):

Thanks guys.